OpenStudy (anonymous):

This is a tutorial Show using the cover-up method that \[ \frac{1}{(x-1)^2 \left(x^2+1\right) (x+1)}=\frac{x+1}{4 \left(x^2+1\right)}-\frac{3}{8 (x-1)}+\frac{1}{4 (x-1)^2}+\frac{1}{8 (x+1)} \]

OpenStudy (anonymous):

\[\frac{1}{(x-1)^2 \left(x^2+1\right) (x+1)}=\frac{A x+B}{ \left(x^2+1\right)}+\frac{C}{ (x-1)}+\frac{D}{ (x-1)^2}+\frac{E}{ (x+1)} \] Let us find D Multiply the two sides by \((x-1)^2\) and let x=1 we get \[ \frac{1}{4}=D \] Multiply the two sides by \(x+1\) and let x=-1 We get \[ \frac{1}{8}=E \] Multiply the two sides by \(x^2+1\) and let x=i We get \[ \frac{1}{(i-1)^2 (i+1)}= A i + B\\ \frac{1}{ }+\frac{i}{4} = A i + B\\ A=B=\frac 1 4 \] Let x=0, we get \[ 1= B -C+ D +E\\ C= B+D+E-1=\frac 1 4+ \frac 1 4 +\frac 1 8-1=-\frac 3 8 \]

OpenStudy (anonymous):

Small misprint for finding A and B \[ \frac{1}{(i-1)^2 (i+1)}= A i + B\\ \frac{1}{4 }+\frac{i}{4} = A i + B\\ A=B=\frac 1 4 \]

OpenStudy (anonymous):

very nice point letting x=i ; i've never seen it thank u sir...

OpenStudy (anonymous):

informative and well-written, great work @eliassaab

OpenStudy (anonymous):

Thank you guys.

OpenStudy (anonymous):

As a practice show using the cover-up method that \[ \frac{100}{(x-1) (x+1)^2 \left(x^2+4\right)}=\frac{4x -4} {x^2+4}-\frac{9}{x+1}-\frac{10}{(x+1)^2}+\frac{5}{x-1} \]

mathslover (mathslover):

gr8 tutorial thanks